Prove that if (G,*) is a finite group, and g is an element of G, then there exists a positive integer n such that g^n =e
1
Expert's answer
2020-10-14T17:04:58-0400

Consider the sequence of powers of an element "g\in G":

"g, g^2, g^3,\dots, g^k,\dots"

Since "G" is finite and "G" contains this sequence, there exist "t, s\in\mathbb N, t<s", such that "g^s=g^t". Multiply both part of this equality by "(g^t)^{-1}=g^{-t}". Then "g^{s-t}=e" and "s-t>0". Let "n=s-t\in\mathbb N". We conclude that "g^n=e."